Bill Rapaport’s

GUIDE TO RESTAURANTS IN AND AROUND BUFFALO

Incorporating Judy and Alan Duchan’s Restaurant Guide

Issue 10, No. 9 — 16 August 1995

                http://www.cs.buffalo.edu/pub/WWW/restaurant.guide/restaurant.guide.html

The Buffalo area has a large number of very good restaurants. I have listed (1) my favorites; (2) those on or near either the UB Amherst or the UB Main St. campuses; and (3) some others recommended by friends, having special cuisines, or otherwise noteworthy (e.g., a perfect or near-perfect rating by The Buffalo News’s Janice Okun).

The list has two parts: §1 is a listing by name only, organized into categories by cuisine and by location. The categories are not intended to be exclusive or exhaustive; corrections to my attempts at categorization are welcomed. §2 is an alphabetical listing of all the restaurants, with details. The selection of restaurants listed here and the comments are my own or those of readers [indicated by initials in brackets]. You should not construe absence from this list as a negative evaluation.

Reservations are always recommended (the restaurant business is volatile; your favorite restaurant might have closed since your last visit!). Call ahead for precise hours of operation; also, some restaurants close for part of the summer.

By the way, the area has two excellent gourmet/ethnic food stores: Premier Center (supermarket-sized, with an adjoining wine and liquor store of equal size; 3455 Delaware Ave., Kenmore; 877–3574) and Tops International (the largest supermarket in, I think, New York State; near the UB Amherst campus, at Maple Rd. and N. Bailey Ave., Amherst; 834–2955).

Some Geography: The UB North (a.k.a. Amherst) Campus is in the Town of Amherst, a suburb northeast of the City of Buffalo. Also in Amherst are the communities of Getzville (northeast of campus) and Eggertsville, Snyder, and the Village of Williamsville (south and east of campus). The Town of Clarence is a suburb east of Amherst. The Town of Tonawanda, separated from Amherst by Niagara Falls Boulevard, is west of campus; the City of Kenmore is in Tonawanda. The Town of Cheektowaga is a suburb south of campus and east of Buffalo; the airport is here. The Village of Depew is in Cheektowaga, southeast of the airport. The UB South (a.k.a. Main Street) Campus is in Buffalo proper, at the northeast corner of the city, about 3 miles southwest of the Amherst Campus (hence its name). Also in Buffalo is Buff State (i.e., SUNY College at Buffalo, not to be confused with UB), on Elmwood Avenue across from the Albright–Knox Art Gallery, west of Delaware Park (one of Olmstead’s creations, by the way, and the location of the Buffalo Zoo). “Allentown” is a historic district of antiques shops, art galleries, mansions, and Victorian houses, about 2 miles south of Buff State and 1 mile north of downtown Buffalo. E. Aurora, Eden, Hamburg, Lackawanna, and Orchard Park are suburbs south of Buffalo. All of the above are in Erie County. Outside of Erie County: Fredonia (home of SUNY Fredonia), is about an hour’s drive southwest of Buffalo on the Thruway. Lewiston (home of Artpark) and Niagara Falls (home of Niagara Falls, of course) are about 25 minutes north of the UB Amherst Campus. Youngstown is north of Buffalo, about 15 minutes from Artpark. In Canada: Fort Erie, Ontario, is across the Peace Bridge, about 10 minutes from downtown. St. Catherine’s, Ontario, is about a 45-minute drive from the Peace Bridge. Niagara-on-the-Lake, Ontario (home of the Shaw Festival) is about an hour’s drive north of Buffalo; the prettiest way to get there (or to Niagara Falls, for that matter) is to cross the Peace Bridge and drive up the Niagara Parkway. Ridgeway is west of Ft. Erie.

Thanks for suggestions to:

[AK] Adam L. Klein  [DK] Deepak Kumar  [JS] Jack Sanders  [MYW] Min–You Wu
[ALS] Alan L. Selman  [DM] Davin Milun  [J&AD] Judy & Alan Duchan [PF] Patsy Fox
[AL] Alan Lockwood  [DN] Debashish Niyogi  [KM] Karin Michelson  [PJE] Patricia J. Eberlein
[AN] Anna Nenonen  [DS] D. Sivakumar  [KNL] K. Nicholas Leibovic  [RH] Robin Hill
[AR] Anthony Ralston  [DWS] David W. Shucard [KR] Kenneth Regan  [RJM] Robert J. Makar
[AS] Annalisa Sassano  [GD] George DeTitta  [KS] Keith Sanders  [RM] Russ Miller
[ATC] An–Tzu Chin  [HD] Harry DeLano  [LH] Lynne Hewitt  [RW] Richard Wyatt
[BJ] Bharat Jayaraman  [JB] Jasti K. Babu  [MLS] Morton L. Schagrin  [SCS] Stuart C. Shapiro
[BS] Barry Smith  [JCM] Jeffrey C. Murphy [MO] Mitsunori Ogiwara  [SC] Sreejit Chakravarty
[CC] Christopher J. Crowner [JH] Jon Hull  [MR] Mary Rapaport  [SF] Scott Field
[CH] Christopher Habel  [JL] Joe Lammens  [MT] Mariam Thalos  [SNS] Sargur N. Srihari

I hope this list contributes to the enjoyment of your stay in the area. Please send comments, corrections, and suggestions to me:

William J. Rapaport

Department of Computer Science

State University of New York at Buffalo Buffalo, NY 14260

     rapaport@cs.buffalo.edu

1 THE RESTAURANTS BY CATEGORIES

The restaurants below are listed by non-exclusive and non-exhaustive categories; within categories, they are listed alphabetically, with my recommendations first (in rank order, alphabetically within ranks). Some names have been abbreviated. Suggestions for improvement of these categories are solicited. For an explanation of location codes [a+], [a], [m], see §2. My rating-code [*]s mean roughly the following:

none = I haven’t eaten here (yet) or I have and wasn’t impressed.
[*] = I’ve eaten here at least once and liked it.
[**] = I’ve eaten here at least once and really liked it.
[***] = Can vie with the best of restaurants in major cities, in my humble opinion.

1. My rankings differ from those of others, and in some cases only reflect my not having eaten there. So a restaurant with 0 stars might be worthy of [***] stars. Read the comments before deciding not to eat somewhere!

2. The UB Commons is a mini-mall on the UB Amherst Campus that has several fast-food restaurants (listed below under “On UB Amherst Campus”), shops and services (including a pharmacy-less drug store and a small grocery store), and offices.

3. “Near UB Amherst Campus” means within 2 miles as the crow flies; except for the fast-food restaurants at the UB Commons, nothing is within walking distance, unfortunately (with the possible exception of the Audubon Room).

4. Buffalo chicken wings are a local specialty that have gained widespread popularity; the restaurants listed below under “Chicken Wings” do them well.

1.1 BY CUISINE

Amish  Chinese  Continental  Dessert  French
Amish Essenhaus [**] Golden Buffalo Dynasty [***] Just Pasta  [***] Just Pasta  [**] Rue Franklin
[**] May Jen/Elmwood  [***] Olivers  [***] Ristorante Giardino  Daniel’s
Cajun  [m**] May Jen/Univ. Plz.  [***] Prince of Wales  [**] Bijou  Michel’s
[**] Rita’s Crystal Palace  [***] White Inn  [a(+?)*] Audubon
Shepherd’s  [a*] Chang’s Garden  [a**] Café in the Square [*] Camp(o)bello’s  Greek
[am*] Mandarin Buffet  [a**] Fanny’s  [a*] Jenny’s
California  [*] Ming Teh  [**] Queen’s Landing  [a*] Johnny B’s  [am**] Family Tree
Bamboo China  [*] Café Cosmos  [a*] Original Pancake House [**] Towne
[**] Bijou  [m] Chinatown  [*] Coda  [m*] Parkside  Ambrosia
[a*] ZuZon  Lee’s  [a*] Port of Entry  [*] Soda Bar  [am] Cobti’s
[m] King’s Wok  [a*] Spina’s  [*] Sweet Tooth  Kostas
Caribbean  [a] New York, NY  Buttery  [m] Ben & JJerry’s
Saki’s  [m] Cameo’s  Butternut Inn  Indian
Babalu  [am] Szechuan House  [a] Daffodil’s  [a] Dessert Deli
Curly’s  [a+] Young Chow  Park Lane  [a] Perkins  [**] Tandoori
Shepherd’s  [am*] Indian Clay-Oven
Chicken Wings (see also: French, English [am*] Raj Mahal
German, Italian, Swiss,  [a] Ashoka
[***] Anchor Bar & Not Otherwise Reginald’s [am] Taste of India
[am*] Duff’s  Categorized)

[a] Elmo’s

[am] Santora’s

Italian  Pizza  Vegetarian  Chains/Franchises
[***] Pranzo  [***] Gino’s  [***] Just Pasta  [am*] Boston Mkt.
[***] Ristorante Giardino [***] Scotto’s  [am*] Juicery  [a*] Original Pancake House
[**] Bflo Rome  [am*] Bocce Club  [m] N. Buff. Food Co-Op [am*] Pizzeria Uno
[**] Little Talia  [am*] Pizzeria Uno  Preservation Hall  [a*] Swiss Chalet
[a**] Spag. West.  [a*] Rigoletto’s  Rutabaga’s  [m] Ben & Jerry’s
[a*] Angelina’s  [a*] Siena  (see also:  [am] Bennigan’s
[*] Camp(o)bello’s  Bijou  Mexican, Chinese,  [a+] Burger King
[m*] Fiorentino  [am] Leonardi’s  Indian, Italian,  [am] Chi–Chi’s
[am*] Pizzeria Uno  [a+] Pizza Hut  Japanese, Pizza,  [a] Country Kitchen
[a*] Rigoletto’s  [a] Pizza Plant  Thai, Vietnamese)  [am] Four Seasons
[*] San Marco  [a] La Porta’s  [am] Friday’s
[a*] Siena  Prima Pizza Pasta  Vietnamese  [am] Fuddrucker’s
Anchor Bar  Raymondo’s  Haxn Haus
Bella Sicilia  [am] Santora’s  [m] Da Tha’o  [a] Perkins
Biac’s  [am] Saigon Bangkok  [a+] Pizza Hut
Café Gar Angelo  Polish  [a] Ponderosa
Coat of Arms  [a] Red Lobster
Left Bank [*] Polish Villa [a] Scotch & Sirloin
[a] Mama Rose’s  Polish Villa II
[a] Olive Garden  Polonia
Prima Pizza Pasta  Strictly Polish
Romanello’s  (and various
Salvatore’s  restaurants
Spaghetti Warehouse  in the
Vito’s Mkt.  Broadway Mkt.)
Japanese  Puerto Rican
Jiro  El Fogon
Saki’s  Niagara Café

Shogun

Seafood

Korean

[**] Harbour Inn

Lee’s  [a**] Rigoletto’s
[m] Seoul House  Captain’s Cove
[a] Woo Chon Coat of Arms Not Otherwise Categorized

Crawdaddy’s

Kosher-Style  Hourglass  [***] Albright-Knox  Fiddle Heads
Jerry’s Seafood  [a(+?)*] Audubon  Flying Tigers
[*?] Stumpy’s  Kodiak  [a*] Deli Bar  Gabriel’s Gate
J.D.’s NY Deli  [a] Red Lobster  [a*] Johnny B’s  Garcia’s Irish Pub
Mastman’s  [am*] Max Hart’s  Justine’s
Swiss  [a*] SoupHerb Gourmet Hemingway’s
Lebanese  Asa Ransom  Hutch’s
Truffles  Billy Ogden’s  Keystone Kelly’s
[am] Cobti’s  Butternut Inn  Lily Langtry’s
Thai  Café in the Promenade  Lord Chumley’s
Mexican/Southwestern  [a] Caffe Aroma  Oban Inn
[a**] Jasmine Thai  Calumet  Olde Fort Inn
[m**] El Charro  [**] Jasmine Thai II Cole’s  [m] Parkside Candy
[a**] Spag. West.  Chon Buri  Country Breads  Pautler’s Drive-In
[a*] Mexican Joe’s  [m] Da Tha’o  [a] Creek View  Pettibone’s
Garcia’s Mexican  [am] Saigon Bangkok Curve  Premier Café
Gramma Mora’s  Saki’s  Cybele’s  Roycroft Inn
[am] Chi–Chi’s  E.B. Green’s  [a+] Sub Shoppe
Lone Star Fajita  [a] Eagle House  [a+] That Coffee Place

[a+] Uni-Mart Warren’s

1.2 BY LOCATION

[a+] On UB Amherst Campus  [a] “Near” UB Amherst Campus  [m] “Near” UB Main St. Campus

(at the UB Commons)

[**] Café in the Square  Ashoka  [**] Café in the Square
Burger King  [**] Family Tree  Bennigan’s  [**] El Charro
Pizza Hut  [**] Fanny’s  Caffe Aroma  [**] Family Tree
Sub Shoppe  [**] Jasmine Thai  Chi–Chi’s  [**] Fanny’s
That Coffee Place  [**] Spag. West.  Cobti’s  [**] May Jen
Uni–Mart  [*] Angelina’s  Country Kitchen  [*] Bocce Club
Young Chow  [*] Audubon  Creek View  [*] Boston Mkt.
[*] Bocce Club  Daffodil’s  [*] Duff’s
Hertel Ave. [h]/  [*] Boston Mkt.  Dessert Deli  [*] Fiorentino
Elsewhere in Bflo  [*] Chang’s Garden  Eagle House  [*] Indian Clay–Oven
[*] Deli Bar  Elmo’s  [*] Johnny B’s
[h**] Little Talia  [*] Duff’s  Four Seasons  [*] Juicery
Billy Ogden’s  [*] Indian Clay–Oven  Friday’s  [*] Mandarin Buffet
[h] Café Gar Angelo  [*] Jenny’s  Fuddrucker’s  [*] Max Hart’s
El Fogon  [*] Johnny B’s  Leonardi’s Pizza  [*] Parkside Candy
Gramma Mora’s  [*] Juicery  Mama Rose’s  [*] Raj Mahal
[h] J.D.’s NY Deli  [*] Mandarin Buffet  New York, NY  [*] Siena
[h] Jerry’s Seafood  [*] Max Hart’s  Olive Garden  Ben & Jerry’s
[h] Kostas  [*] Mexican Joe’s  La Porta’s Pizzeria Bennigan’s
[h] Lone Star Fajita  [*] Original Pancake House Perkins  Cameo
[h] Mastman’s  [*] Pizzeria Uno  Pizza Plant  Chi–Chi’s
Niagara Café  [*] Port of Entry  Ponderosa  Chinatown
[*] Raj Mahal  Red Lobster  Cobti’s
Downtown Bflo  [*] Rigoletto’s  Saigon Bangkok  Da Tha’o
[*] Siena  Santora’s  Four Seasons
[**] Bijou  [*] SoupHerb Gourmet  Scotch & Sirloin  Friday’s
[*] Juicery  [*] Spina’s  Szechuan House  Fuddrucker’s
[*] Swiss Chalet  [*] Swiss Chalet  Taste of India  Hourglass
Calumet  [*] ZuZon  Woo Chon Korean King’s Wok
Crawdaddy’s  Leonardi’s Pizzeria
Curve  West Side:  N. Buff. Food Co-Op
E.B. Green’s  Allentown/Elmwood/Buff State areas  Saigon Bangkok
Garcia’s Irish Pub  Seoul House
Hemingway’s  [***] Albright–Knox  Biac’s  Szechuan House
Justine’s  [***/wings] Anchor Bar  Cole’s  Taste of India
Keystone Kelly’s  [***] Just Pasta  Cybele’s
Pettibone’s  [***] Oliver’s  Fiddle Heads
Prima Pizza Pasta  [**] Bflo Rome  Gabriel’s Gate
Saki’s  [**] May Jen  Hutch’s
Spaghetti Warehouse  [**] Rue Franklin  Kodiak
Vito’s Market  [**] Tandoori  Left Bank
[**] Towne  Lord Chumley’s
[*] Café Cosmos  Park Lane
[*] Coda  Preservation Hall
[*] Juicery  Reginald’s
[*] Soda Bar  Romanello’s
[*] Sweet Tooth  Rutabaga’s
Ambrosia  Taste of India
Babalu  Vito’s Mkt.

Suburbs: North & East = Suburbs: South =  Outdoor Dining (Summer)
(Other) Amherst/ E. Aurora/Eden/ (probably incomplete list)
Cheektowaga/Clarence  Hamburg/Lackawanna/
(Other) Tonawanda Orchard Park [***] Just Pasta

[**] Bflo Rome

[***] Pranzo  Amish Essenhaus  [**] Bijou
[***] Scotto’s Pizza  Butternut Inn  [am**] Café in the Square
[**] Golden Buffalo Dynasty Country Breds  [m**] May Jen/Univ. Plz.
[**] Jasmine Thai II  Curly’s  [a**] Jenny’s
[**] Rita’s Crystal Palace  Daniel’s  [a*] Johnny B’s
[*] Boston Mkt.  Polonia  [a*] SoupHerb Gourmet
[*] Camp(o)bello’s  Roycroft Inn  [*] Sweet Tooth
[*] Pizzeria Uno  Calumet
[*] Polish Villa Outside Erie County Crawdaddy’s
[*] San Marco  Hemingway’s
[*?] Stumpy’s  [***] Gino’s Pizza  Reginald’s
[*] Swiss Chalet  [***] Prince of Wales  Vito’s Mkt.
Asa Ransom  [***] Ristorante Giardino (and UB Commons)
Bamboo China  [***] White Inn
Bella Sicilia  [**] Harbour Inn
Café on the Promenade  [**] Queen’s Landing
Captain’s Cove  [*] Ming Teh
Flying Tigers  Buttery
Garcia’s Mexican  Chon Buri
Jiro  Coat of Arms
Haxn Haus  Michel’s
Kostas  Oban Inn
Lily Langtry’s  Olde Fort Inn

Olive Garden Pautler’s Drive-In Polish Villa II Premier Café Raymondo’s Salvatore’s Shepherd’s Shogun Tandoori Truffles Warren’s

2 ALPHABETICAL LISTING OF RESTAURANTS

Key: [*], [**], [***] = Bill’s favorites  [XXX] = recommendation from person listed above whose initials are XXX
[NEW] = newly listed in this volume  [a+] = on UB Amherst Campus (in the UB Commons)
[REV] = revised from previous volume [a] = near UB Amherst Campus (within 2 miles as the crow flies)

[m] = near UB Main St. Campus (within 2 miles as the crow flies)

[***] Albright–Knox Art Gallery. 1285 Elmwood Ave. (in Delaware Park, across from Buff State), Buffalo. 882–8700. Sun. brunch (reservations required) and excellent lunches in a beautiful setting. Run by the same people who run Café in the Square. The Gallery’s collections span the full gamut of art; Smithsonian Magazine says it has the best collection of late-20th C. art in the U.S. [RM, 5/94] says: “Fun to eat lunch and tour the gallery. Food is reasonably priced and OK. Don’t expect a gourmet meal.”

[REV] Ambrosia. 467 Elmwood Ave. (just north of Allentown), Buffalo. 881–2196. [J&AD, 1993] say: “A Greek-American restaurant offering standard Greek clichés, well prepared. Family owned, friendly service. Located in a neighborhood which is interesting to explore.” [AN&BS, 5/94] say: “Good cheap Greek place (better than, e.g., Towne ’cos does non-soggy food). All the snack things and starters are good, and service fast and frendly even if you are in a hurry.”

Amish Essenhaus. 8151 N. Main St., Eden. 992–9544. Amish cuisine in a converted farmhouse next to the Eden Cheese Factory.

[REV ***] Anchor Bar. 1047 Main St. (at North St., near Allentown), Buffalo. 886–8920. The place that invented Buffalo chicken wings and, in my opinion, have the best ones. Their Italian food is OK (of the spaghetti-and-meatballs variety), but people come here primarily for wings. [J&AD, 1993] say: “We recommend it so you will know what the authentic taste [of wings] is. The other food is OK, but not memorable.” [RM, 5/95] says: “Real good wings—a Buffalo tradition—this is where it all started.”

[REV a*] Angelina’s Italian Gourmet. 952 Maple Rd. (east of N. Forest, in Boardwalk Boutiques Plaza), Williamsville. 689–3011. [BJ, 6/94] says: “Another good nearby Italian restaurant, like Rigoletto’s. Simple but elegant decor.” [AN, 7/95] says: “Mediocre.”

Asa Ransom House. 10529 Main St., Clarence. 759–2315. [RM, 5/94] says: “There is a big fuss about this and I don’t understand why. It is a nice old place, and fun to see, but the food itself is just average and the prices a bit high.” (The “big fuss” may, for all I know, be due to the former chef, who left for the White Inn several years ago.) Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[NEW a] Ashoka Indian Cuisine. 3246 Sheridan Dr. (in Hills Plaza, near Bailey), Amherst. 831–8608. This is the latest restaurant in what [SCS] calls the “Indian Restaurant Generator”, since there have been a series of Indian restaurants that open up here and then move elsewhere.

[REV a(+?)*] Audubon Room. 2401 N. Forest Rd. (in the University Inn), Amherst. 636–7500. Probably the closest restaurant off the UB Amherst Campus; walkable in nice weather, maybe. I’ve had several nice lunches here. But a dinner I had (1993) was rather mixed; I had a very nice swordfish with fruit salsa, but [MR]’s chicken tasted as though it had been cooked in the morning and then nuked just before being served. [MR] and I agree that desserts were exquisite, however. [RM, 5/95] says: “Absolutely top shelf. One of the best 5 or so restaurants in the Buffalo area. The lamb is wonderful as are the pasta dishes. Listen carefully to the specials, and if something strikes your fancy, order it. Lunches are also really interesting, well-presented, and reasonably priced. A real hidden treasure.” Given 4 \frac{1}{2} (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV] Babalu. 41 Virginia Place (Allentown), Buffalo. 882–2822. Caribbean. [AK, 5/94] reports: “Jamaican cuisine; good to very good quality; service unknown (I’ve never been—my sister loves it).” [AN, 7/95] says: “Good food, interesting flavours.”

[REV] Bamboo China. 6600 Main St. (between Transit and Youngs), Williamsville. 633–5033. [JH, 6/94] says that it “seem[s] to be [one of] the best Chinese restaurants in town. Bamboo China is really pretty good. Especially if you order one of their specialties. They are the only place we’ve been able to find that has good noodle soup. On a cold wet Buffalo day, a bowl of their beef noodle soup really hits the spot. There are a few other good dishes at Bamboo China, too (one of them is something like ‘Four treasures … ’).” Based on this recommendation, [MR] and I ate here; we hated it. [AN, 7/95] says: “I don’t like it much, either.”

[REV] La Bella Sicilia. 2909 Genesee St. (east of Harlem Rd.), Cheektowaga. 896–6692. Italian. [DWS, 5/95] says that he was taken here “when we first came to Buffalo and it is still excellent. They have great linguine and red clam sauce and their pizza is excellent. I still eat there. Also their antipasto is great.”

[m] Ben & Jerry’s Ice Cream. 59 Kenmore Ave. (at Windermere, near University Plaza, west of UB Main St. Campus), Amherst. 834–8566. Famous for their Vermont ice cream. [DK]

[am] Bennigan’s. 4060 Maple Rd., Amherst. 838–3550. Part of the chain.

[REV **] Bflo Rome Café. 291 Bryant St. (west of Elmwood Ave., just north of Allentown), Buffalo. 881–6747. Not as in “Oh, give me a home where the”; rather, mostly northern Italian cuisine. Everything is à la carte. Very small, so be sure to make reservations. [J&AD, 1993] say: “This tiny place, with a Greenwich Village atmosphere, offers an eclectic menu, with interesting, well-presented dishes. Many of the dishes come in both small and large portions, saving on cost and calories.” [GD, 1993] says: “Hands down the best place on Bryant—makes Just Pasta look like just what it is—a yuppie refuge.” [AR, 3/94] says: “Mediocre meal there last Sunday evening.” [AK, 5/94] says: “Small (seats ˜15–20 tops); somewhat expensive (entree ˜ $7.00–$15.00); food of consistent very good to excellent quality; gourmet cuisine; wonderful Sunday brunch (get the stuffed French toast); very good to excellent service.” [AN, 7/95] says: “Some dishes are very good, some not what you’d want them to be.” WARNING: No non-smoking section. Outdoor dining in summer.

[REV] Biac’s World Bistro. 581 Delaware Ave. (in Allentown; near Theodore Roosevelt inauguration site), Buffalo. 884–6595. Mediterranean/Italian. I, Bill, ate here in summer 1993. I had a good but not outstanding corn soup; an excellent arugula, tomato, and cucumber salad; and a grilled tuna special that reeked of smoke flavor (which I despise); [MR] had an antipasto, which she didn’t like. Unbearably noisy. WARNING: The non-smoking section changes from day to day, according to one waiter, and we were seated—by the owner, no less—in the smoking section even though we explicitly asked for non-smoking. [ALS, 6/94] says: “Highly recommended. Biac’s is only unbearably noisy on Friday nights. Other nights are fine. Of course, if you are looking for the Yuppie bar scene on a Friday night, then maybe Biac’s is where you want to be.” [PJE, 5/95] says: “Biac’s is noisy and not very good food either.” [AN, 7/95] says: “Actually very nice, except for the geriatric yuppies.” Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[**] Bijou Grille. 643 Main St. (downtown, next to Market Arcade Cinema, across from Theater Place), Buffalo. 847–1512. Self-described as “California-style bistro dining.” Originally from the folks who brought you Just Pasta (below), now under new management, but otherwise unchanged. Informal, nice place to eat downtown. [J&AD, 1993] say: “This 50s’ style popular downtown nightspot features sandwiches, salads, a few entrees, and flavorful single-serving pizzas.” Also, good desserts; wine by the glass. Outdoor dining in summer (downtown Buffalo really looks good from here).

[REV] Billy Ogden’s Lovejoy Grill. 1834 William St. (at Ogden [get it? “William (Billy) Ogden”?], in south Buffalo), Buffalo. 896–8018. [AN, 7/95] says: “Nice seafood.” Given 3 \frac{1}{2} (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[NEW am*] Boston Market. 4185 Maple Rd. (next to Pizzeria Uno, across from the University Cinema), Amherst. 834–2320. Also at 8160 Transit Rd. (in East View Plaza, north of Maple Rd.), Williamsville; 688–1946. Part of the chain.

[a+] Burger King. UB Commons. 633–8700. Part of the chain.

Butternut Inn. Rt. 16 & Genessee St., Chaffee (east of East Aurora). 496–8987. [KR, 6/94] says: “Nice food in nice atmosphere, but IMHO the soups and bread and vegetables weren’t as fresh or as interesting as in other comparable restaurants. Dessert, however, was fully up to Just Pasta standards, with a home-baked country flair.”

[REV] The Buttery. 19 Queen St., Niagara-on-the-Lake, Ontario, Canada. 905–468–2464. [SCS] has said: “Directly across the street from the Court House Theater; à la carte entree $15–$20 (Canadian); my favorite in Niagara-on-the-Lake; eating outside on the porch in good weather is very pleasant; also have ‘Henry VIII Feaste’ available.”, but he also says (5/95) that although he hasn’t eaten there in several years, he still has fond thoughts of it. On the other hand, I [Bill] ate here twice a long time ago and didn’t like it either time.

[REV *] Café Cosmos. 206 Allen St. (west of Elmwood Ave., in Allentown), Buffalo. 885–0502. Warning: Rumor has it that they are out of business, so call before going! From a chef formerly at Just Pasta, Calumet, and Biac’s. [MR] and I ate here for the first time in [10/94]; I was generally pleased with my meal (a rather bland mushroom soup, a laughably miniscule salad, very good trout in a walnut butter sauce, and a nice pear tart for dessert) (however, the service, though attentive, was a bit on the slow side, worrisome since we had a concert to go to that night and we kept eyeing the clock). [MR] was less pleased (bland chicken, herbed mashed potatoes that kept repeating during the concert). I’d certainly be willing to try it again. [ALS, 6/94] says: “Cosmos is excellent and very creative.” [JS, 6/94] says: “A very fine restaurant.”

[NEW] Café Gar Angelo. 1197 Hertel Ave., Buffalo. 875–5057. [AL, 7/95] says: “I have eaten twice at Café Gar Angelo, an Italian restauraunt on Hertel that I think is the best food deal in Buffalo. Complete meals for less than $10, fairly good wine for less than $15 per bottle; I’ve liked everything I had there, but thought the veal Gar Angelo was terrific.”

[am**] Café in the Square. 4476 Main St. (Snyder Square Shopping Plaza), Snyder. 839–5330. Excellent gourmet lunches and dinners. Run by the same people who run the restaurant in the Albright–Knox. Outdoor dining in summer. Jazz and dinner theater on weekends. The best chocolate-chip cookies around.

[NEW] Café on the Promenade. 4243 Genesee St. (in the Radisson Hotel, across from the airport), Cheektowaga. 634–2300. [RM, 5/95] says: “Brunch at the Radisson is one of the best around town. Lots of choices, decent atmosphere, and the food is very well prepared. It appears they have recently changed bakeries, and the desserts are suffering a bit, though they are still pretty good.”

[NEW a] Caffe Aroma. 5229 Main St., Williamsville. 631–2687. Coffee, pastries, Panini sandwiches on foccacia. [AN, 7/95] says: “Good luncheons.”

[REV] Calumet Arts Café. 56 W. Chippewa St. (downtown), Buffalo. 855–2220. Boasts a “performance room” for exhibits, readings, etc. [RH, 9/93] says: “Had Saturday lunch there, which was fine, but quite standard. Met one of the owners, Mark Goldman, who inquired about my French fries, tried one, pronounced them awful, and promised better ones, dropping dire hints about a new cook. I think they are betting their reputation on the Sunday music brunch, which should be worth a try.” [J&AD, 1993] say: “The Calumet offers well-prepared, unusual, artfully presented lunches and dinners. The setting is congenial and historic. They offer jazz or folk music on weekends, with a reasonable cover charge.” [AN, 7/95] says: “Food doesn’t live up to the rest of it (music, decor, etc.).”

[REV m] Cameo’s. 482 Niagara Falls Blvd. (near Kenmore Ave.), Tonawanda. 832–2790. Continental cuisine. A few years ago,

[DK] said: “The food here is consistently good!” [MYW, 6/95] recommends it, too.

[REV **] Camp(o)bello’s Café. 9416 Transit Rd. (north of Casey, in Amherst Place Plaza), E. Amherst. 689–1141. Good Italian cuisine. Informal. Huge portions. Good desserts. However, [RM, 5/94] says: “Has continued to be somewhat disappointing, though it is always crowded, has reasonably priced meals, and is not terrible for lunch. Certainly is not a true café.” Note: Despite what it says on the entrance or in the phone book, the name is really ‘Campobello’s’. [MYW, 6/95] says: “Good food, inexpensive. Great linguine with red clam sauce. Cakes [come] from Sweet Tooth.” They also have a pizza place around the corner at 9334 Transit Rd. (at Casey, in Transit–Casey Plaza), E. Amherst, 636-8770, about which [RM, 5/95] says: “Nice Italian pizzeria. Pretty decent food at a very nice price. I have had most of their menu items. Everything but the pizza is pretty good.”

[NEW] Captain’s Cove. 829 Dick Rd., Cheektowaga. 632–4810. [MYW, 6/95] says: “One of the best seafood restaurants. Fresh, well-prepared seafood. Inexpensive.”

[REV a*] Chang’s Garden. 935 Maple Rd. (east of N. Forest, in Boardwalk Boutiques Plaza), Amherst. 689–3355. Has a new “northern China” menu (which I haven’t tried yet) in addition to their regular fare, but my most recent meal here (7/95) was quite good—far better, in fact, than my most recent meal at Rita’s Crystal Palace, which I generally prefer to Chang’s. Perhaps they have a new chef (see [RM]’s comment, below). [AN&BS, 5/94]’s favorite Chinese restaurant (“best dish: General Tso’s Chicken”). [BJ, 6/94] says: “Fairly good vegetarian food. I enjoyed their garlic eggplant.” [JH, 6/94] says: “In my opinion, it’s one of the best in Buffalo.” Also highly recommended by [SNS]. And [MYW] said (and still [5/95] agrees): “It is a good Chinese restaurant in the Buffalo area. They did pretty nice work in food decoration. The food itself tasted good, although they put a little bit more sugar in every dish. And I recommend the Peking duck.” [RM, 7/95] says: “While this is quite popular around town, I have never been taken with the food. On a recent visit, however, I had some of the more exotic dishes, and found them quite tasty and well prepared.” [DM, 7/95] “much prefers” Chang’s Garden to Ming Teh. Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV m**] El Charro. 3447 Bailey Ave. (at Winspear, south of UB Main St. Campus), Buffalo. 837–5300. Very good Mexican food; not a chain like Chi–Chi’s. They serve the best fajitas in town. [KR, 6/94] says: “Good and steady; lighter fare than Mexican Joe’s. Pleasant dining room and good service.” [AN, 7/95] says: “My favourite Mexican” restaurant in Buffalo.

[am] Chi–Chi’s. 4003 Maple Rd., Amherst. 834–2281. Mexican. Part of the chain.

[m] Chinatown. 51 University Plaza (across from UB Main St. Campus), Buffalo. 834–6615. [DN, 1993] says: “Mainly take-out Chinese food, although they do have a few tables if you want to eat in. Very good & inexpensive food, large portions. Free Chinese tea to sip while you watch your food being prepared.”

Chon Buri. 208 Church St., St. Catherine’s, Ontario, Canada. 416–687–8424. [DN, 1993] said of this place when it was in Welland, Ontario: “Exquisite Thai food, very spicy.” Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV] Coat of Arms. 425 Center St., Lewiston. 754–2311. A few years ago, [MLS] said: “A small, inexpensive Italian restaurant. Superb pasta dishes along with interesting fish choices (not necessarily Italian). Pickled olives served with drinks; home baked bread; 3 house white wines. Reservations probably a good idea on busy weekends.” He still (5/95) agrees with that, though he notes he hasn’t been there in a while.

[am] Cobti’s. 1927 Eggert Rd., Eggertsville. 833–6967. Greek and Lebanese.

[*] Coda. 350 Pennsylvania (on the West Side, across the street from the Saarinen-designed Kleinhans Music Hall, where the Buffalo Philharmonic plays), Buffalo. 886–6647. A very small restaurant, beautifully decorated, and well-situated for those attending a concert at Kleinhans or a play at the Kavinoky theater. The menu is also very small: 2 appetizers, 3 entrees, 3 desserts the time I went (the menu changes regularly). The food is not superb, but it is interestingly prepared [10/93].

[NEW] Cole’s. 1104 Elmwood Ave. (between Allentown & Buff State), Buffalo. 886–1449. [JCM, 5/95] says: “Good sandwiches and entrees; standard menu of mixed meat and fish dinners. Many different styles of sandwiches. Large dining room with booths for multiple people and a smaller intimate section for couples. Good bar, with lively bar-tenders, off of the main dining room. Prices are average for Buffalo. Informal.”

Country Breads & More. 1089 Davis Rd. (Rt. 240), West Falls (near Orchard Park). 655–0039. [RJM, 5/95] says: “My wife and I love the place . . .. You might take the 219 and get off at Armour Duells to get there. Its a great place to go as part of a ride in the country on a Saturday or Sunday. The place is part bakery and part restaurant. They are only open from 9 A.M. to 4 P.M. They serve great breakfasts and lunches. There are dinners on the menu, but we haven’t tried them. They bring a nice basket of baked goods to your table while you wait for your order. The place can get crowded, and you might have to wait for your table and your order, but the food has always been worth it. My favorites are the Country Wife omelet and the meat pies. We always take home bags of cookies and breads when we leave.”

[NEW a*] Country Kitchen. 800 Maple Rd., Williamsville. 688–1232. Standard, but good, local chain of coffee shops. Formerly a Four Seasons (see below). [RM, 5/95] says of this reincarnation: “A major improvement, but still a mediocre diner.”

Crawdaddy’s. 2 Templeton Terr. (downtown, overlooking Lake Erie), Buffalo. 856–9191. OK seafood; absolutely fantastic view of waterfront, marina, skyline of downtown Buffalo. [J&A, 1993] say: “Great views of the Buffalo marina and decent seafood if you stick to the simple presentations.” Once upon a time, this place was given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer, but I doubt that rating is valid any more.

[a] Creek View. 5629 Main St., Williamsville. 632–9373. [AN&BS, 5/94] say: “Decent Yank/Irish food; good value.” Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV] Curly’s Bar and Grill. 647 Ridge Rd., Lackawanna. 824–9716. Caribbean, Asian, and continental. [DN, 6/94] says: “Their Jamaican Jerk Chicken is excellent. Also be sure to try their delicious Conch Fritters appetizers.”

The Curve. 132 Genessee St. (downtown), Buffalo. 852–5075. [J&AD, 1993] say: “This is a new restaurant to Buffalo. It is well decorated, featuring paintings from local artists. The meals are attractive and light. The breakfasts are especially appealing.”

[REV] Cybele’s. 99 Elmwood Ave. (Allentown), Buffalo. 882–1247. [RH, 9/93] says: “Sue and I, having failed to make it to the Grad Picnic in time for the food, found ourselves looking for a place to eat on Elmwood, Sunday afternoon at 3:30—a bad time. We ended up at Cybele’s, just north of Allen Street, where the staff was very accommodating (but not exactly quick) in spite of the fact that they were recovering from serving a large poetry reading. The food is quite good, in the modern Bflo Rome style. Portions were small, but so were prices.” [AN, 7/95] says: “Nice atmosphere, if you like cats.”

[REV a] Daffodil’s. 930 Maple Rd., Williamsville. 688–5413. I have eaten here twice and not liked it either time: mediocre and pretentious food. If you go, take a flashlight so you can see where you are. However, [RM, 5/94] says: “One of the best restaurants in Williamsville. The lamb chops are a specialty and worth the price. Interesting atmosphere if you enjoy a library-type setting.” More recently, [JS, 10/94] says: “We tried Daffodil’s the other day in spite of your pan, because I was really hungry for lamb; a quick scan of the “Guide” came up with only Daffodil’s mentioning lamb as a specialty. They weren’t serving lamb chops this time, but they did have the rack of lamb, which the waiter billed as “world-famous”. Service was friendly—even jovial, as in the case of the possibly only half-serious comments on world-famous rack of lamb and carrot cake. I got the half-rack at $23, and it was nicely prepared. I wanted it pink and got it that way. It was served with a very nice sauce which (unfortunately) left me wondering whether a bit of A-1 had crept in. My younger son … got a “small” slice of prime rib (he doesn’t have much of an appetite for such things), which was still rather large and quite good. In the end, I’d recommend it. But I would not call it a world-class experience.” And [KS, 10/94] says: “The lemon pepper chicken was a bit excessively lemony, but still quite enjoyable. Salad and dressing alike were fresh and tasty. The carrot cake was world-famous. The atmosphere was definitely upscale, but we weren’t made to feel uncomfortable in our sport coats and jeans.”

[REV] Daniels. 174 Buffalo St., Hamburg. 648–6554. French. Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[m] Da Tha’o Vietnamese and Thai Restaurant. 3106 Main St., Buffalo. 838–0670.

[NEW a*] Deli Bar. 5542 Main St. (between N. Cayuga and Rock, near the pastoral Glen Park (worth a visit for its waterfall and duck pond)), Williamsville. 633–0733; fax 633-0734. Coffees, soups, sandwiches, desserts. [AN, 7/95] says: “I was disappointed.”

[REV a] Dessert Deli. Maple–Forest Plaza, Williamsville. 689–2115. Take-out bakery. There’s also a [NEW a] Dessert Deli Café in Corvette Cleaner Plaza, around the corner on N. Forest.

[am*] Duff’s. 3651 Sheridan Dr. (near Millersport Hwy.), Amherst. 834–6234. The best wings near the UB Amherst Campus.

[NEW a] Eagle House. 5578 Main St. (near the pastoral Glen Park (worth a visit for its waterfall and duck pond)), Williamsville. 632–7669. [RM, 5/95] says: “Nice atmosphere; entertainment in the lounge on some weekend nights. Decent food in a nice, cozy atmosphere. A favorite of people who live in the village of Williamsville.” However, I [Bill] think the place is pretty mediocre.

[REV] E. B. Green’s. 2 Fountain Plaza (in the Hyatt Regency Hotel, downtown), Buffalo. 856–1539. Steakhouse. [DM, 7/95] says: “Very good steak—all the way up to the 48 oz. porterhouse!” Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[NEW a] Elmo’s. 2349 Millersport Hwy., Getzville. 688–7237. [JB, 5/94] says: “Their Cajun wings are the best and their wings can beat Duff’s any day. The prices are affordable (5 lbs. of wings (30 of ’em) cost 10 bucks). We the folks at the Computing Center demand that it be put up among the best of the Buffalo’s best. You should try it, too.”

[am**] Family Tree. 4346 Bailey Ave. (at Ruth, about midway between UB Main St. & Amherst Campuses), Eggertsville. 838–2233. Greek. Formerly a branch of the Towne (see below), now a solo operation. Has a larger menu (e.g., more seafood, wider variety of desserts) than the Towne, but is otherwise indistinguishable.

[REV am**] Fanny’s. 3500 Sheridan Dr. (east of Millersport Hwy.), Amherst. 834–8665. Elegant, expensive, continental cuisine. Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[NEW] Fiddle Heads. 62 Allen St. (in Allentown), Buffalo. 883–4166. [RM, 7/95] says: “Nice new restaurant. Very interesting, well-prepared, and moderately priced lunches. They are now crowded for dinner, so make reservations. If you are in the Allentown area, definitely drop in for lunch.” Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV m*] Il Fiorentino. 727 Kenmore Ave. (1 block west of Starin), Kenmore. 833–4175. Italian. [J&AD, 1993] say: “Leisurely dining in a chef-owned family-run Tuscan restaurant. Expensive, and worth it.” [ALS] thinks that this is one of the 3 best restaurants in the area, saying “I really do love Il Fiorentino, not only for the food but for the charm and ambiance as well”. [AR, 9/93] says: “It certainly is expensive by Buffalo standards but not, I think, worth it. Reasonably good but not worth more than your one star. As to the ambiance, I found it more pretentious than charming.” [AN&BS, 5/94] say: “Best restaurant we’ve found in Buffalo. [It is] maybe even better than Saki’s for a romantic night out.” [PJE, 5/95] says that meals “take too long” here. Given 4 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer. [BS, 7/95] says: “Still probably the only real restaurant in WNY. Why does this [expletive deleted by editor :-)] Rapaport give Il Fiorentino only one star?” (Editor’s reply: Because I’ve only eaten here once, and it was a special meal, not the regular menu.)

[NEW] Flying Tigers. 100 Amherst Villa Rd., Cheektowaga. 631–3465. [MYW, 6/95] says: “A restaurant owned by the China-based ‘Flying Tiger’ team in World War II. Nice atmosphere. Oversees runway of the airport. Food is fine.”

El Fogon. 235 Niagara St. (near Carolina), Buffalo. 845–5137. [PF, 1994] says: “This is one of two restaurants serving standard but very good Puerto Rican food.”

[REV am*] Four Seasons. 3388 Sheridan Dr. (at Sweet Home Rd.), Amherst, 837–0702. Standard, but good, chain of coffee shops.

[am] Friday’s. 1209 Niagara Falls Blvd. (in Boulevard Mall), Amherst. 835–6000. Part of the chain.

[am] Fuddrucker’s. 4300 Maple Rd. (near Sweet Home Rd.), Amherst. 836–3683. Part of the chain.

[NEW] Gabriel’s Gate. 145 Allen St. (east of Elmwood Ave., in Allentown), Buffalo. 886–0602. [RM, 7/95] says: “Nice, relaxing lunch. When the patio is open, that is also pleasant. Standard, fairly large, lunch and dinner menu.” I [Bill] have been here a few times; it’s OK, but nothing special. However, [DM, 7/95] says that it “has some of the best wings in Buffalo.”

[NEW] Garcia’s Mexican Restaurant. 3755 Union Rd., Cheektowaga. 681–9595. [MT, 5/95] says: “Tex/Mex cuisine; the traditional stuff is ordinary, but their tuna rubbed in chile, with salsa and vegetables, is outstanding. Virgin drinks from the bar (called smoothies) are out of this world.”

Garcia’s Irish Pub. 74 Pearl St. (downtown), Buffalo. 856–0111. [J&AD, 1993] say: “A popular singles’ bar, Garcia’s offers sandwiches, salads, and light entrees.”

[REV ***] Gino’s New York Pizza. Rainbow Mall, Niagara Falls. One of the two best pizzerias in Western New York! It really tastes like New York pizza!! [CC, 5/95] agrees: “… the pizza place in the Rainbow Mall, which I thank you for identifying in your restaurant guide!, makes me feel … like I’m in NYC (culinarily) …;-).”

[**] Golden Buffalo Dynasty. French and Union Rds. (in Garden Village Plaza), Cheektowaga. 656–9053. Off the beaten path, but a good Chinese restaurant (indeed, one of my two current favorites in the area, the other being Rita’s Crystal Palace). Offers Peking Duck every day, without a reservation. I also liked their Moo Shu Chicken and their Lovers in the Stars (beef with scallion pancakes).

[REV] Gramma Mora’s. 2051 Niagara St. (at Hertel, on the west side), Buffalo. 877–9532. Some [HD, 5/95] say that it has good, inexpensive, fairly genuine Mexican food, with a “very homey” atmosphere; others say that it has bad, inexpensive, non-genuine Mexican food :-). [J&AD, 1993] say: “An inexpensive popular place . . .. The meals are plentiful, atmosphere is fun, but hectic.” [DN, 1993] says: “The most authentic Mexican food in Buffalo.”

[REV **] Harbour Inn. 35 Melville St., Niagara-on-the-Lake, Ontario, Canada. 905–468–3881. Excellent seafood.

[NEW] Haxn Haus. 1840 Maple Rd. (at Ayer, in Maple–Ayer Plaza), Williamsville. 639–9020. [RM, 8/94] says: “Really nice, pleasant, family place. Rotisserie cooking. I have had the chicken and beef, and enjoyed both. Will certainly go back and try some of their other dishes. Their sauce is quite tasty. Cafeteria style.”

Hemingway’s. 492 Pearl St. (downtown), Buffalo. 852–1937. [J&AD, 1993] say: “Standard American food at this bar and restaurant. Check to see if jazz is offered for the night you want to go.”

[m] Hourglass. 981 Kenmore Ave. (near Colvin Ave.), Kenmore. 877–8788. [J&AD, 1993] say: “A small, family owned restaurant. Everything is good. The seafood is outstanding. The wine list is extensive and fairly priced.” Given 4 \frac{1}{2} (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[NEW] Hutch’s. 1375 Delaware Ave. (south of Delaware Park), Buffalo. 885–0074. Bistro run by former chef at Oliver’s. [PJE, 5/95] says: “Pretty good.” Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV am*] Indian Clay-Oven. 940 Millersport Highway (at Sheridan Dr., across the street from Duff’s), Amherst. 832–1030. I [Bill] have had several very good meals here in the last year (1994). [LH, 6/94] says: “Clay Oven is generally improved.” [AN, 7/95] says: “Pretty good.”

[REV a**] Jasmine Thai Restaurant. 1330 Niagara Falls Blvd. (across from Boulevard Mall), Tonawanda. 838-3011. I have had many exquisite meals here [3, 5, 7, 9/94, etc., etc.], though it’s not a typical, big-city Thai restaurant. I have never been disappointed, nor have I gotten bored, even after eating here several days in a row during the Summer Institute in Cognitive Science (where we took many visitors, all of whom liked it very much). In contrast, [DN, 1993] says: “Not too bad, but nowhere near the quality of Chon Buri.”

[NEW **] Jasmine II Thai Restaurant. 3719 Union Rd., Cheektowaga. 683–6553. Not a duplicate of its elder sibling (see above)—it has a different, and equally good, menu. Slightly more elegant surroundings than Jasmine I. [AN, 7/95] says: “Decent food.”

[NEW] J.D.’s New York Deli and Restaurant. 1434 Hertel Ave. (near the North Park Theater, which shows foreign films), Buffalo. 836–3313. Kosher-style deli.

[a*] Jenny’s Ice Cream. 78 E. Spring St., Williamsville. 633–2424. Good ice cream, sandwiches, and salads, in a nice setting, near the pastoral Glen Park (worth a visit for its waterfall and duck pond). Outdoor dining in summer. WARNING: No rest room!

Jerry’s Seafood Market and Restaurant. 1226 Hertel Ave., Buffalo. 876–4138. [J&AD, 1993] say: “The name speaks for itself. An offshoot of an established Buffalo fish-market. Very well prepared seafood.” [MLS, 5/95] says: “A small to medium-size place, with spare decor. Wide selection of fresh fish cooked to perfection. I prefer less elaborate sauces, but I am not complaining about the chef’s imagination. Prices slightly above moderate. WARNING: Too small for a non-smoking section. We sat next to a table with two chain-smokers.” Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

Jiro. 8316 Main St. (near Transit Rd.), Williamsville. 632–2229. Japanese; sushi bar. Lunches. [MO, 1992] says: “You can enjoy typical Japanese food with reasonable price. Decoration is not so awful as many Japanese restaurants have. Sushi is fair, but not excellent. The problem might be in their using larger amount of rice than usual. Sukiyaki is bad; it’s too sweet and not with any Japanese vegetables. Deep fried materials like Tonkatsu (pork fry), Tempura (shrimp, pumpkin, and so on) are good.” [AN&BS, 5/94] say: “Halfway decent Sushi bar.”

[am*] Johnny B’s. Maple Ridge Centre (near the Maple Ridge cinema), Maple and Sweet Home Rds., Amherst. 838–1950. Note the phone number: this place tries to remind you of a ’50s soda shop. It’s pretty good. Outdoor dining in summer. [DN, 1993] says: “Fairly good sandwiches, sundaes.” [RM, 5/94] says: “Remains an old fashioned 50s diner, with juke box and all. Nice for burger and fries or club sandwich before a movie. Ice cream is not too bad either.”

[REV am*] Juicery. Century Mall/Northtown Plaza (Sheridan near Bailey), Eggertsville. 833–2360. Cafeteria-style, mostly vegetarian. Cheap and good. [BJ, 6/94] says: “A nice place for a quick meal, especially for vegetarians.” [DN, 1993] says: “Try their pita sandwiches with their delicious Miso dressing.” Other locations in The Bon-Ton department store in Eastern Hills Mall, 4545 Transit Rd. (north of Main St.), Williamsville, 634–4020; in Delaware Park, 873–3035; and downtown, at 110 Theatre Place, Buffalo, 855–1216.

[REV] Justine’s. 120 Church St. (in the Buffalo Hilton, downtown, on the waterfront), Buffalo. 845–5100. Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV ***] Just Pasta. 307 Bryant St. (west of Elmwood Ave., just north of Allentown), Buffalo. 881–1888. My favorite. Menus change with the season. Not just pasta, but also all sorts of other good things with pasta as a side dish. The desserts here are extraordinary; my favorite is White Satin. Informal; not too expensive. Good list of wines by the glass. Good lunches, too. And outdoor dining in summer. [BJ, 6/94] says: “I visited this place with a faculty candidate in 1990,* and they improvised an excellent vegetarian primavera for me. My visit in 1993 was also with a faculty candidate, but this time I was a bit disappointed with the food.” [PJE, 5/95] says: “Just Pasta is expanding and has had spirits for quite a while now. I’ve had super meals there. They’re always crowded.” “A tier down” from one of [KNL]’s “top-ranked favorites” (5/95). Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer. (* Note: [BJ] informs me that both faculty candidates are now on our faculty; Just Pasta wins again!)

Keystone Kelly’s. 101 Theater Pl. (downtown), Buffalo. 854–1900. [J&AD, 1993] say: “Campy, campy, campy. A hangout for young singles. Decent sandwiches and light food.”

[m] King’s Wok. 3233 Bailey Ave. (south of UB Main St. Campus), Buffalo. 837–1740. Good Chinese. Szechuan. [MS] says: “King’s Wok did indeed start with excellent food and decline a bit, but its food is still excellent, particularly for the prices; still among the best Chinese restaurants in the area (Ming Teh is the best).” Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV] Kodiak Café. 423 Elmwood Ave. (near Bryant, just north of Allentown), Buffalo. 886–4226. Seafood. [PJE, 5/95] says: “OK, but pricy for what you get. I had a strange combination of pea soup with salmon, and the peas were al dente(?).”

[REV] Kostas. 9210 Transit Rd. (south of Casey), E. Amherst. 639–8009. Greek. Similar sort of Greek menu to Towne and Family Tree, but they also have taramosalata (fish roe dip), melitzanosalata (eggplant salad), saganaki (broiled Kaseri cheese), and tzatziki (yogurt, cucumber, and dill salad), which are really good (though the melitzanosalata is a bit bland for my tastes) [11/93]. [RM, 5/95] says: “Pretty nice new diner. They have a wide selection and it is quite spacious and airy inside. Food is decent.” Also: 1561 Hertel Ave. (east of North Park Theater, which shows foreign films), Buffalo. 836–9944.

[REV] Lee’s. 290 Oliver St., North Tonawanda. 693–5241. [JHL, 5/91] says: “They serve Korean and Chinese dishes and they are pretty good.” More recently, [HD, 7/95] says: “It’s run by a Korean family. They have Korean and Chinese food. Korean food is delicious and nicely served. Real homey kind of place.”

[REV] Left Bank. 511 Rhode Island St. (near Richmond; on the West Side, near Allentown), Buffalo. 882–3509. [RM, 5/94] says: “Very nice pub-like atmosphere. Neighborhood-type of friendly crowd. Casual and relaxing. The food is moderately priced and pretty good, though one needs an apetite for garlic, as that is the mainstay of most of the Italian-style dishes.” [AN, 7/95] says: “Great ambience.”

[NEW am] Leonardi’s Pizzeria. 614 Grover Cleveland Hwy. (near Eggert and Longmeadow). 835–8700. [DS, 1/95] says: “Their white pizza (no tomato sauce) is excellent; comparable to or better than Bocce’s. (Maybe even [Bill’s] ‘NYC taste buds’ will like it!).”

[NEW] Lily Langtry’s. 6261 Transit Rd., E. Amherst. 688–4746. [RM, 5/95] says: “Nice atmosphere and moderate prices. The food is fair.”

[**] Little Talia Trattoria. 1458 Hertel Ave. (near North Park Theater, which shows foreign films), Buffalo. 833–8667. [AS] (a native Italian) says: “Genuine and sophisticated Italian food. Friendly service, elegant interior (charming and authentic family pictures on the wall—with unfortunate spelling errors in the captions!!). Family run by Italian-Americans.” [JL] (a native Belgian) says: “The best Italian food I’ve tasted in Buffalo. Home cooked and fresh tasting. A small place, and they don’t take reservations. Moderately priced (we paid $45 for two).” The best seat in the house is the booth for two in the non-smoking section upstairs that overlooks the kitchen; you can watch a cooking show as entertainment while you dine!

[REV] Lone Star Fajita Grill. 1853 Hertel Ave. (between Parkhurst and Starin), Buffalo. 833–7756. [PF] says: “They serve fajitas, tacos, chile, Tex–Mex fashion, and have lots of magazines about Texas lying around. It’s a déclassé but hip atmosphere.” [KR, 6/94] says: “Takeout or no-frills sit-down. Inexpensive Tex–Mex with all-fresh ingredients and a distinctive flair. The polar opposite of Chi–Chi’s. Very simple and tasty. If you’re into ‘slumming’ (a pre-PC ’70s term—the correct usage rules out McDonald’s etc., and would be translated as ‘roving for authenticity’ today), this rates 5 stars.” [AN, 7/95] says: “ ‘Cool’.”

Lord Chumley’s. 481 Delaware Ave. (Allentown), Buffalo. 886–2220. [J&AD, 1993] say: “Great Caesar salad, elaborate verbal menu, congenial service, intimate atmosphere.” [AK, 5/94] says: “Still very good food; somewhat expensive; excellent service; very good Caesar salad.” Given 3 \frac{1}{2} (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV a] Mama Rose’s. 2555 Millersport Hwy., Getzville. 688–1555. Basic Italian restaurant. [RM] says: “Small, nice atmosphere; reservations recommended near weekends; pretty good basic Italian food. Inexpensive to moderate price.” [RM, 5/94] says: “Unpretentious Italian restaurant. I like their sauce and they have a wide range of tasty meals that can satisfy all ranges of groups. Nice, quiet, relaxing atmosphere. Good salads.” [MYW, 6/95] recommends it, too.

[NEW am*] Mandarin Buffet. 60 Maple Rd., Amherst. 633–6222. [MR] and I have been here twice, once for lunch (quite good; excellent hot-and-sour soup), and once for dinner (not so good—maybe a different chef?). However, it was far superior to the only other Chinese buffet we’ve tried (King Wah, on Sheridan, in Williamsville Place). Note that it has non-Chinese dishes, too, including a very good salad-and-fruit bar and a good dessert bar. [MYW, 6/95] says: “Best Chinese buffet in town. Huge selections.” [RM, 7/95] says: “Wide selection of Chinese and American foods. Lunch is much better than dinner and fairly enjoyable. I have been several times, and it has not been terribly busy. I wonder what will next occupy this location.” (That’s a reference to the fact that the restaurants that have preceded it at this location have all gone out of business.)

Mastman’s Kosher Delicatessen. 1322 Hertel Ave., Buffalo. 876–7580. Kosher-style.

[**] May Jen/Elmwood. 810 Elmwood Ave. (between Allentown & Buff State), Buffalo. 881–0038. Cantonese & Szechuan. I used to think this was far inferior to its less elegant sibling at University Plaza, but my last meal here was far better than my last meal at University Plaza; maybe they’ve switched chefs. [J&AD, 1993] say: “In our opinion, the best Chinese restaurant in Buffalo (but not as good as Fort Erie’s Ming Teh).” [AR, 9/93] says that he had “a good meal”.

[m**] May Jen/University Plaza. 47 Kenmore Ave. (near UB Main St. Campus), Buffalo. 832–5162. Cantonese & Szechuan. I used to think this was far superior to its more elegant sibling in the Allentown/Elmwood Ave. area, but my last meal here was not as good as my last meal at the Elmwood Ave. location; maybe they’ve switched chefs. Outdoor dining in summer. WARNING: no non-smoking section; however, they have expanded since my last visit, so this warning may no longer be needed.

[REV am*] Max Hart’s Diner. 20 Flint Rd., Amherst. 689–2211. Sun–Thurs, 7 A.M.–11 P.M.; Fri–Sat, 7 A.M.–2 A.M. Between the Hampton House and Red Roof motels; not otherwise noteworthy, though I have had decent lunches here (9/94).

[a*] Mexican Joe’s. Maple and N. Forest Rds., Williamsville. 632–4530. Not quite as good as El Charro, but far superior to Chi–Chi’s.

Michel’s. 3589 Garrison Rd., Ridgeway, Ontario, Canada. 416–894–4444. French. Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV *] Ming Teh. 126 Niagara Blvd., Fort Erie, Ontario, Canada. 416–871–7971. Szechuan. [MS] says: “Ming Teh is the best.” Considered by many to be not only the best Chinese restaurant in the area but of world-class quality. You can request no MSG; reservations a must [MLS]. However, I [Bill] must admit I was a bit disappointed with my last meal here. [J&AD, 1993] say: “This Chinese restaurant is just across the border. It offers unusual dishes, well presented, good service, and has a nice view overlooking the Niagara River.” [J&AD, 1993] think that Ming Teh is better than May Jen/Elmwood, which they think is the best Chinese restaurant in Buffalo. [GD, 1993] says: “There are no other Chinese restaurants worth the time while Ming is on the block.” But [MLS] says: “I concur that Ming Teh has slipped. It is still good, but not really worth the fuss. I’d just as soon go to May Jen.” [RM, 5/94] says: “Had been the best Chinese in the area for some time. Still the best, but it is noticeably slipping.” [DM, 7/95] says: “Went there for the first time recently, after hearing much about their ‘wonderful’ food. Went with a large group. I was very disappointed. A couple of the dishes were unusual and tasty, but most were very average. I much prefer Chang’s Garden.” Given 4 \frac{1}{2} (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[a] New York, NY. 1376 Niagara Falls Blvd. (at Maple), Tonawanda. 832–4491. Despite its name, this was one of the best Chinese restaurants in the area when I ate there the first time (but why do they serve a lettuce salad with each meal?). But it has deteriorated; the last meal I had here was the last meal I will have here. [RW] says: “The first time I was there, which was in the second week after its opening, we all thought it was truly excellent; since then, I’ve been there four more times, and each time it was (at least, seemed) worse than the previous time; the four of us there last week all agreed that it was sufficiently lousy to not consider going there again.” [SF, 11/93] recommends them “for their Dim Sum brunch which they have every Sat. & Sun. morning starting about 10:30. It’s as close to NY or Toronto as you’re going to get. Carts with assorted dumplings are wheeled around and you can pick and choose. There are many choices from the standard shiu mai and war tip to items such as chicken feet and tripe. They also have a grill cart for fried items. (Stay away from the sate though, the pork was tough.) The selection seems to vary from week to week.”

[REV] Niagara Café. 525 Niagara St., Buffalo. 885–2233. [PF] says: “This is one of two restaurants serving standard but very good Puerto Rican food.” [CC, 5/95] says: “A good place for Latin-American food—more specifically Puerto Rican/Cuban-style food, not Mexican or Caribbean (the kind you can get in thousands of places in the NYC area): ropa vieja, carne asada, bistec, black bean soup, yellow rice, etc. Only [Gino’s] pizza place in the Rainbow Mall … makes me feel as much like I’m in NYC (culinarily) as this place

;-).”

[NEW m] North Buffalo Food Co-Op Vegetarian Café. 3144 Main St., Buffalo. 836–8058. [HD, 5/95] says: “Delicious and varied selection, big portions at reasonable prices, and very nutritious.”

[NEW] Oban Inn. 160 Front St., Niagara-on-the-Lake, Ontario, Canada. 905–468–2165, 905–468–7811. One of [KNL]’s “top-ranked favorites” (5/95).

Olde Fort Inn. 110 Main St., Youngstown. 745–7141. Near Old Fort Niagara, about 15 minutes from Artpark. Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV a] Olive Garden. 3951 Maple Rd. (next to Scotch and Sirloin, near Bailey), Amherst. 837–5211. Also: 2011 Walden Ave. (near Walden Galleria Mall), Cheektowaga. 683–4670. Part of the chain. Reviews are for the (older) Walden Ave. location. [RM, 5/94] says: “The food has been mediocre at best each of the times I have been.” [BJ, 6/94] says: “A moderately good, but not so nearby, Italian restaurant. The (vegetarian) waitress who served us mentioned that the tomato sauce used in most of their entrees contains beef fat.”

[REV ***] Oliver’s. 2095 Delaware Ave. (north of Delaware Park), Buffalo. 877–9662. Elegant, expensive, continental cuisine. I prefer Oliver’s to Fanny’s (but Fanny’s is no slouch, either). [AR] thinks that Oliver’s is good, but not at all in the same league as Rue Franklin; [AR, 8 Sep 94] said: “First meal there in some time last Sunday. Only average. Good appetizer, salad; mediocre entree.” [J&AD, 1993] say “Elegant, expensive, innovative, excellent.” However, [AN] says: “Oliver’s has gone downhill. All the starters we (a party of three) had were soggy, salty, and horrible; even the oysters had been killed beyond recognition by cheese and onions. The service was downright rude. The meat was excellent. But stop recommending this place!” I [Bill] went there shortly after reading [AN]’s comments, and found the place as good as ever, so either our tastes differ or they were there on an off-night. However, I was there on a Tuesday evening, and it was unbearably noisy. [BS, 6/94] reiterates [AN]’s comments: “I haven’t been to Oliver’s since; it was horrible. But maybe the waiter has been fired. However, I would not go there again; I’d go to Il Fiorentino, which maybe is even better than Saki’s for a romantic night out.” WARNING: They allow smoking in the non-smoking section. Given 4 \frac{1}{2} (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[a*] Original Pancake House. 5479 Main St., Williamsville. 634–5515. For those old-timers who still remember La Crèpe or Le Piano à Crèpes, yes, there still is a restaurant in the area that serves crèpes! If the following list sounds good, try this place; it is good: kijafa cherry crèpes, crèpes suzette, mandarin orange crèpes, peach crèpes, mushroom crèpes, beef stroganoff crèpes, chicken crèpes, seafood crèpes; also: apple pancake, German pancake. Huge portions. Also serves standard coffee-shop fare. Extremely crowded on weekends, especially Sunday brunch-time, but fast service once you get in. They do not have any smoking section!

Park Lane. Delaware Ave. at Gates Circle (between Allentown and Buff State), Buffalo. 883–3344. [DN, 6/94] says: “Beautifully presented food, attentive but unobtrusive service. Varied and tasteful continental cuisine. Very elegant, old-world decor and ambiance (the suggested dress code is ‘nice casual’, but don’t let that fool you).”

[m*] Parkside Candy Co. 3208 Main St. (at Winspear, west of UB Main St. Campus), Buffalo. 833–7540. Ice cream, (homemade) candy, good sandwiches. A fantastic interior.

[NEW] Pautler’s Drive-In. 6343 Transit Rd., E. Amherst. 688–6463. [RM, 5/95] says: “Nice place to go with the kids for a hot dog and/or ice cream after a baseball or soccer game.”

[a] Perkins. 3920 Maple Rd., Amherst. 837–0296. Pancakes, etc. Part of the chain.

Pettibone’s Grille. NorthAmericare Park, 275 Washington St. (downtown), Buffalo. 846–2110. NorthAmericare Park is the newly renamed baseball diamond of the Buffalo Bisons. [J&AD, 1993] say: “Pettibone’s Grille is located in the stadium, and overlooks the field. It serves standard American food, nicely prepared.”

[a+] Pizza Hut. UB Commons. 689–6507. Part of the chain. [RM, 5/94] says: “Poor pizza, but a reasonable salad bar.”

[REV am] Pizza Plant. 3085 Sheridan Dr. (Northtown Plaza), Amherst. 833–0882. Also, [a] 5110 Main St. (Walker Center plaza), Williamsville; 626–5566. Also, 8020 Transit Rd., Williamsville. 632–0800. Stuffed pizza (“pods”). Vegetarian pizzas available. [BJ, 6/94] adds: “A good place for exotic pizzas. Vegetarians would find quite a selection of inexpensive foods.” On the other hand: My [Bill’s] New York City taste buds cringe at this place, but locals seem to like it. [RM, 5/95] says: “Specializing in calzones, which they like to call pods. A truly confusing menu, which was cute years ago when it was a bit smaller. While I have never enjoyed their food, many native Buffalonians like it (though they like Bocce’s pizza as well).” [DM, 7/95] says: “(Since there are many negative reviews, I think I’ll add a positive one!): I eat here often. I prefer their stuffed pizzas to their pods, but their food is very tasty, and you can have an unimaginably huge selection of toppings. Also has a large selection of beers, many imported.”

[NEW am*] Pizzeria Uno. 4125 Maple Rd. (across from the University Cinema), Amherst. 834–6200. Also in the Walden Galleria Mall, Cheektowaga, 683–0925. Part of the chain.

[*] Polish Villa. 2954 Union Rd., Cheektowaga. 683–9460. Some Polish cuisine. One of our colloquium speakers, who comes from Warsaw, was pleased with the Polish food here. One warning: The “Polish Platter”, with golambki (stuffed cabbage), sauerkraut and cheese pierogi, kielbasa (sausage), and potatoes, is enough to feed 2 people!

[REV] Polish Villa II. 1085 Harlem Rd., Cheektowaga. 882–4908. [MYW, 6/95] says: “A Polish restaurant offering standard Polish food.”

Polonia. 3071 Abbott Rd., Orchard Park. 828–1410. Southern Polish food: pierogi, gołubki, naleśniki, bigos, leniwe, potato pancakes.

[a] Ponderosa Steak House. 4350 Maple Rd. (near Sweet Home Rd.), Amherst. 838–3182. Part of the chain.

[REV a*] Port of Entry. 270 Campbell Blvd., Getzville. 689–7374. [ATC] says: “Located on a parcel of land consisting of a few restored farmhouses which have been converted into country craft stores and a nice, quiet restaurant. Craft stores worth browsing, too. A quiet country setting.” [DN, 1993] says: “Excellent food, nice atmosphere.” [RM, 5/95] says: “Nice restaurant, with lots of neat gift items in the lobby and store behind the restaurant. Good food, moderately priced, nice atmosphere. They are best known for their ‘Oscars’, which I quite like.”

[REV a] La Porta’s Pizzeria. 485 W. Klein Rd. (near Hopkins), Williamsville. 688–7740. [RM, 5/95] says: “The pizza I like best around the Amherst area—they also make a nice antipasto.”

[REV ***] Pranzo. 4243 Genesee St. (in the Radisson Hotel, across from the airport), Cheektowaga. 634–2300. Italian. Superb meals with superb service—an elegant dining experience, much to my surprise given its location. [BJ, 6/94] says: “There wasn’t a single vegetarian entree but they were glad to improvise, and are very good at it. You will be seated in a corner if you are not dressed well. Overall, an excellent but expensive place.” [AN&BS, 10/94] say: “We went to Pranzo finally. ’Twas quite good, really (we both think so). P.S. from [AN]: I won’t comment on the decor because the food was so good. … Although next time I go there I will bring a [flashlight] with me.” [RM, 5/95] says: “Northern Italian at its best. Get dressed up, bring a lot of cash, and enjoy a fabulous meal. Make sure to get a selection of appetizers and have some dessert.” And a few days later, [RM] added: “I was at Pranzo again a couple of nights ago—superb as usual.” Given 4 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer. [MYW, 6/95] recommends it, too.

[NEW] Premier Café. 3455 Delaware Ave. (south of Sheridan), Kenmore. 877–3574. In the wonderful Premier Liquor and Gourmet food center. Eat in or take out. Breakfast, lunch, early dinner. Espresso bar. Sandwiches, salads, entrees, pizzas, cheeses, desserts.

[REV] Preservation Hall Bakery & Café. 532 Elmwood Ave. (just north of Allentown), Buffalo. 884–4242. Recently moved to new, bigger location. Wide variety of vegetarian dishes and soups: stuffed pita sandwiches, pastries, bagels, and homemade preserves. [BJ, 6/94] says: “Another good place for inexpensive vegetarian meals.”

Prima Pizza Pasta. 396 Pearl St. (downtown), Buffalo. 852–5555. [J&AD, 1993] say: “Fast food pizza and other Italian standards. Filling, dependable, and inexpensive.”

[REV ***] Prince of Wales. 6 Picton St., Niagara-on-the-Lake, Ontario, Canada. 905–468–3246. Elegant, moderately expensive, continental cuisine. However, [AR] reports that he “had a quite mediocre meal there a few weeks ago. Not even worth one star.” (But if you’ve read this far, you know that [AR] and I hardly ever agree on restaurants :-)

[REV **] Queen’s Landing. 155 Byron, Niagara-on-the-Lake, Ontario, Canada. 905–468–2195. Elegant, moderately expensive, continental cuisine.

[am*] Raj Mahal. 850 Niagara Falls Blvd. (near Eggert, south of Sheridan), Tonawanda. 832–1588. A Buffalo branch of a Rochester restaurant. A slightly wider menu than Indian Clay-Oven; indeed, at least two Indian friends of mine like Raj Mahal better than Clay-Oven. [SC, 9/93] says: “I think Raj Mahal is also good” although he prefers Taste of India. [BJ, 6/94] says: “In August 1993, there was a change in management and chef, with noticeable effect on the quality of their food. I rate their (vegetarian) food now as Very Good.”

[REV] Raymondo’s Pizzeria. 5651 Transit Rd., Clarence. [RM, 5/95] says: “They make a nice veggie pizza, but the rest of their fare is fairly uninteresting.”

[a*] Red Carpet. Changed its name from ‘Spina’s Red Carpet’ to Spina’s (see below).

[a] Red Lobster. 4010 Maple Rd. (at Bailey), Amherst. 837–0202. Part of the chain.

[REV] Reginald’s. 718 Elmwood Ave. (near Ferry, between Allentown and Buff State), Buffalo. 885–3370. Advertises itself as English cuisine. They also serve Tea (with scones and all that) in the afternoon. [J&AD, 1993] say “This friendly place is informal. It has eclectic continental dishes including some well-prepared British Empire standbys such as meat pies. Service is attentive.” [HD, 7/95] says: “The vegetarian fare has pretty much disappeared.”

[REV a*] Rigoletto’s. Two equally good branches: Maple Ridge Centre (near the Maple Ridge cinema), Maple and Sweet Home Rds., Amherst; 833–4607. Also at 716 Elmwood Ave. (near Ferry, between Allentown and Buff State), Buffalo; 884–0461. Gourmet Italian with an operatic theme; gourmet pizzas. More red sauces than at Just Pasta, but nothing whatsoever like your corner spaghettiand-meatballs place. They also serve matzah-ball soup! [J&AD, 1993] say “The offerings include pasta, salads, and seafood. The atmosphere is informal with a neon-Italian decor. Moderate prices.” [RM, 5/94] says: “I don’t know what the fuss is about. I don’t care for their sauce and have tried [it] several times. Always disappointed.” [BJ, 6/94] says: “I once took a colloquium speaker …, and he greatly enjoyed one of their seafood specials.”

[REV ***] Ristorante Giardino. 142 Queen St. (Gate House Hotel), Niagara-on-the-Lake, Ontario. 905–468–3263. Exquisite, northern Italian food; desserts as beautiful to look at as to eat. Very expensive. Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV **] Rita’s Crystal Palace. 9380 Transit Rd. (north of Casey), E. Amherst. 639–0031. Currently, one of my two favorite Chinese restaurants in the area (the other is Golden Buffalo Dynasty). A large menu with many items not found elsewhere (e.g., hot pots, jellyfish). Their broccoli in garlic sauce is heavenly. [RM, 5/95] says: “If you stick with the non-exotic, it is one of the better Chinese restaurants around. They deliver.”

Romanello’s Roseland. 490 Rhode Island St. (on the West Side, near Allentown), Buffalo. 882–3328. Italian. [MLS] said: “2 dinner specials (soup to after-dinner drink); not expensive. [AR, 3/94] says: “First time there in ages; mediocre would be an overstatement.”

[NEW] Roycroft Inn. 40 S. Grove St., E. Aurora. 652–5552. Recently renovated historic inn, originally built in 1905 as part of Elbert Hubbard’s Arts and Crafts Community.

[REV **] Rue Franklin West. 341 Franklin St. (between Tupper St. and Edward St., near Allentown), Buffalo. 852–4416. Elegant, expensive, French cuisine. Reservations a must. Highly recommended by [AR], who prefers it to Oliver’s; indeed, [AR] and I have a friendly dispute about this place. The last time I went there, our table wasn’t ready, they ran out of some food items (it was only 7:30 P.M.!), and the service was poor. But the food was great. [J&AD, 1993] say: “Buffalo’s only pure French restaurant. For Buffalo, it is expensive, but well worth it.” Sigh, again I must dispute with colleagues: My Parisian friends who ate there hardly considered it as “pure French”. [ALS] thinks that this is one of the 3 best restaurants in the area. [AR, 9/93] says: “You’ll be surprised that I agree with your Parisian friends about ‘pure French’. Although I still think it’s the best restaurant around, I try to avoid it for a few months after I’ve been to France so as not to make the inevitable invidious comparisons.” And [AR, 3/94] says: “Superb; clearly the best meal I’ve had in Buffalo this academic year; in particular, much superior to a good meal at Olivers a few weeks ago!” And [AR, 5/94] says: “Had a superb meal there last night. Confirms my opinion that it is head and shoulders over any other restaurant in Buffalo not just because of the food but also because the ambience is far superior to any other restaurant.” One of [KNL]’s “top-ranked favorites” (5/95). Given 4 \frac{1}{2} (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[NEW] Rutabaga’s. 177 Hodge St. (west of Elmwood Ave., just north of Allentown), Buffalo. 882–2909. Vegetarian.

[NEW am] Saigon Bangkok. 512 Niagara Falls Blvd. (south of Sheridan Dr.), Tonawanda. 837–2115. Thai and Vietnamese. [DS, 1/95] says: “The place is nice and the food is very good. My wife and I have tried pretty much all the vegetarian dishes they have, and they are all good—their ‘Spicy Rainbow’ is one of the best. Another good news for vegetarians is that they use only vegetable oils in all their preparations! Heartily recommended.” [MT, 5/95] says: “Good spicy soups; not as lavish in their use of herbs as authenticity would demand; inexpensive and large portions.”

[REV] Saki’s. 28 Church St. (downtown, in the beautifully renovated, classic Louis Sullivan skyscraper—the Guaranty Building), Buffalo. 852–0012. Chinese-Japanese-Thai-American. [J&AD, 1993] say: “Saki’s offers attractively prepared food with a Pacific Rim influence. It includes a sushi bar.” [MO] says: “They serve Japanese food with a flavor of Thai cooking. Sushi bar is excellent; fish is fresh, the taste of the sushi rice is well prepared. Moreover, the appetizers are quite good. The price is a bit high; a Sushi plate with several separate orders of sushi, an appetizer, and a bottle of beer cost about $30 including tax and tip.” [AN] says: “The best restaurant in town is Saki’s sushi bar.” [AK, 5/94] says: “Expensive (everything à la carte); very good sushi (as well as ‘Pacific rim’ cuisine; eclectic menu; good service.” [PJE, 5/95] likes “the hot shrimp pasta stuff”. Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

Salvatore’s Italian Gardens. 6461 Transit Rd. (not far from airport), Depew. 683–7990. Italian. Known primarily for its amazingly gaudy decor. [MLS] says: “Once past the decor, a decent meal.” After living in Buffalo for over 10 years, I [Bill] finally ate here. The food and service were quite good. The decor is indeed gaudy, though some of it (especially the Christmas decorations) are not bad—a bit like being inside a New York City department store window. However, although we were seated in a non-smoking section (and the restaurant is huge—certainly large enough to have a large non-smoking section), a table of about 10 people were seated right next to us, in the non-smoking section, and all of them were smoking. I would have complained to the owner and the waiter, but ’twas they who seated them—apparently they were good customers. (To be fair, when I asked the waiter if indeed this was the non-smoking section, he replied, rather drily, “it was”).

[REV *] San Marco Ristorante. 2516 Elmwood Ave. (at Tremont), Kenmore. 874–6711. Moving in August 1995 to 2082 Kensington Ave. (just off the I-290 interchange at Main), Snyder. Northern Italian. [GD, 1993] says: “Forget … [what] you hear from J[anice] O[kun] (who has … [no] discernment …) about Il Florentino being the only ‘authentic’ Italian restaurant in town. San Marco is the real thing (we were just in Italy for three weeks, on our fourth trip there, so when we compare San Marco with true Italian restaurants we are not just jabbering). Frank and Nancy Grimaldi have made the place as warm and inviting as is the food well prepared and generous in taste. The strongest points are the game (venison, wild boar, bacchia à la Romagna), which is just unavailable elsewhere in Buffalo. The pastas are all prepared from the get-go. In most ‘Italian’ places in town, the pasta is half boiled, then given the last two minutes to finish it off. There is no way to retain the proper bite in pasta so cruelly abused. Each sauce is finished in an individual pot or pan. Pasta is on your plate and at your table within a minute of leaving the water. The service is warm and professional (no ‘have-a-nice-day’ or ‘and-how-are-we-this-evening’ diabetic nightmares), with waiters who have actually been trained to open a bottle of wine. The wine list is the most sophisticated grouping of Italian wines outside of a very few top restaurants in Italy. Brunellos abound, as well as Barolos, etc. Weak points: Salads could be more interesting, and desserts are very limited …, but if Nancy’s tiramisu is on the menu, don’t ever have it—if you ever want to be pleased by all the other also-rans in town. The marcipone cheese is the sweetest in town, the lady fingers melt, the cocoa is rich and bitter.” [AR, 9/93] says: “Still very good although perhaps not quite up to the standard before I went on sabbatical. But I’ll be back.” [RM, 5/94] says: “Really wonderful family owned Northern Italian Restaurant. Ask Frank what he recommends on any given day and enjoy!” [PJE, 5/94] says: “Very good (better than Just Pasta now, I think).” [BJ, 6/94] says: “I greatly enjoyed the pasta and salad. Not much for vegetarians here, but whatever they had was very good.” [PJE, 5/95] says: “Very good, though I’ve eaten their menu, I think.” [BS, 7/95] says: “Why does this [expletive deleted by editor :-)] Rapaport give San Marco only one star?” (Editor’s reply: Because I’ve only eaten here once.)

[am] Santora’s Take Out & Delivery. 1402 Flint Rd. (across from Marriott and Red Roof, in Triad Apartment Bldg.), Amherst. 688–3081. Delivers subs, pizzas, wings.

[REV a] Scotch and Sirloin. 3999 Maple Rd. (at Bailey), Amherst. 837–4900. Part of the chain. [RM, 5/95] says: “Pretty good steak and a small salad bar. Nice and relaxing.”

[NEW] Scotto’s Pizza. Eastern Hills Mall Food Court, 4545 Transit Rd., Williamsville. 634–3003. One of the two best pizzerias in Western New York! It really tastes like New York pizza!!

[m] Seoul House. 38 Kenmore Ave. (across from University Plaza, west of UB Main St. Campus), Buffalo. 836–3122. Korean.

Shepherd’s. 8222 Transit Rd. (in East View Plaza, north of Maple Rd.), Williamsville. 689–7802. Cajun and Continental. [DN, 1993] says: “Excellent food, well presented. Superb service.” Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[am*] Siena. 4516 Main St., Amherst. 839–3108. Gourmet pizza and pasta. Outdoor dining in summer. Informal but upscale. Although [MR] and I enjoyed our meal here, [KM] had such a bad experience with both food and service that she will never return. [AR, 9/93] says: “Not much better than mediocre. Won’t go back.” WARNING: no non-smoking section.

[REV] Shogun Ichi. 7590 Transit Rd., Williamsville. 631–8899. Japanese teppan bar restaurant. [JCM, 5/95] says: “Very good food and a great show. A little pricy, but worth it. Informal. Tables seat up to parties of 10. Reservations are recommended, since the restaurant does not have many tables. The dishes were very tasty, and the menu had a good selection.”

[*] Soda Bar & Pastry Co. 929 Elmwood Ave. (between Allentown and Buff State), Buffalo. 885–0056.

[NEW [a*] SoupHerb Gourmet. 5403 Sheridan Dr. (in Williamsville Place Plaza), Williamsville. 631–4822. Gourmet soups, sandwiches, salads, light meals. Mostly lunches, but some early dinners.

[REV] Spaghetti Warehouse. 141 Elm St. (downtown), Buffalo. 856–1350. [DN, 1993] says: “Large, busy restaurant for people with big appetites and a liking for Italian food. Fairly good food, large portions. Interesting decor (including a converted railway dining compartment).” [RM, 7/95] says: “Inexpensive, standard Italian food. Not very tasty, but not too expensive. Popular with a younger crowd.”

[REV a**] Spaghetti Western. 15 New Rd. (near Dodge), E. Amherst. 689–0034. A.k.a. “Spag West”. Italian/Southwestern US cuisine. I’ve had both good dinners and good lunches here. [BJ, 6/94] says: “Their pasta puttanesca is excellent—can be had without anchovies.” [RM, 5/94] says: “A favorite. A mixture of Italian and Tex-Mex cuisine that seems to hit the spot whenever I go. Nice relaxing atmosphere as well. Interesting daily specials.” [CH, 7/94] says: “The tex-mex chicken dishes are great at Spaghetti Western. We think that the ** grading is appropriate.” Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV a*] Sprina’s. 5507 Main St., Williamsville. 634–1968. Formerly, ‘(Spina’s) Red Carpet’. Run by a former chef from Oliver’s (see above), so it’s quite good. But can be very noisy. Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[NEW *?] Stumpy’s Delights. 3347 Delaware Ave. (just south of Premier Center gourmet food store), Kenmore. 873–2010. Formerly Ralph’s on Hertel. I haven’t tried Stumpy’s yet, but here’s what was said of Ralph’s (which I rated *): Jewish-style (but not kosher) deli. WARNING: Ralph’s had no real non-smoking section; only “non-smoking tables” (I don’t know about Stumpy’s). [GD, 1993] said: “Deli at its western New York best. The pastrami and the corned beef are great, but try the brisket—it’s the finest.”

[a+] Sub Shoppe. UB Commons. 639-0694.

[REV *] Sweet Tooth. 478 Elmwood Ave. (just north of Allentown), Buffalo. 884–2520. Also: [a] 15 Rock St. (north of Main, near Cayuga, and near the pastoral Glen Park (worth a visit for its waterfall and duck pond), Williamsville. 632–6232. The Allentown location serves candy, ice cream, sodas, wonderful pastry, bagels, sandwiches. Self-service. Limited seating. Outdoor seating in summer. [MYW, 6/95] says: “Best cake in town. We buy birthday cake only from there. Both [locations] are good. But Allentown has more selections.”

[NEW a*] Swiss Chalet. 1551 Niagara Falls Blvd., Amherst, 837–4300. Also: 5205 Transit Rd., Clarence; 634–6000. Also: 660 Main St. (downtown), Buffalo, 852–0008. Part of the chain. For a chain, it’s not bad; good chicken soup. You can ask for all white meat. [RM, 5/94] says: “Chicken and ribs, made reasonably well and moderately priced. Nice family restaurant that won’t wipe out your wallet.”

[NEW am] Szechuan House. 839 Niagara Falls Blvd. (near Eggert), Amherst. 838–9114. [DS, 5/95] says: “Has very good options for vegetarians. Their Hunan Eggplant, Buddhist Pot, Pan-Fried-Noodles-with-Vegetables, and Vegetable Fried Rice were all very good (the eggplant was great).”

[REV **] Tandoori. 445 Delaware Ave. (Allentown), Buffalo. 847–1112. [NEW] Also on 7200 Transit Rd. (at Main St., in the Econo Lodge Motel), Williamsville; 632–1112. (Recommendations are for the Allentown location.) In my opinion, this is the best Indian restaurant in the area, even though my clothes smelled like curry afterwards! [J&AD, 1993] say “Attractive Indian restaurant with an extensive menu. Owned by members of an extended family who also have restaurants in Jerusalem, London, and New York City.” [SC, 9/93] says: “Tandoori is not even ‘authentic’ in that: What they list is not what they serve. Out of the three items I ordered, two were very bad. It is also very expensive.” [LH, 6/94] says: “For a more upscale experience, I’d also recommend Tandoori—also overall good quality, with fancier decor and higher prices, but worth it, as some dishes are interestingly different from the standard North Indian fare.” [PJE, 5/95] says: “Tandoori has a good new menu, and I can now order there without having Indians with me.” Given 3 (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV am] Taste of India. 3093 Sheridan Dr. (in Northtown Plaza), Amherst. 837–0460. 484 Elmwood Ave. (just north of Allentown), Buffalo. 881–3141. [SC, 9/93] says: “I personally think that the Indian food in Taste of India [Amherst branch] is the best.” Also about the Amherst branch, [LH, 6/94] says: “My recommendation for a modest Indian restaurant is Taste of India; overall good quality, good service, nice buffet lunch; dosas on Sundays not bad (not that I’ve ever had a really good dosa in North America, but still); I’ve never had anything awful there.” [PJE, 5/95] says: “Had a good lunch with [an Indian colleague] at the Taste of India in Northtown Plaza. It was a ‘buffet’ and really very good. But [my colleague] says it’s better sometimes than others. I suspect that Thurs–Fri has fresher food than Mon–Tues.” And [HD, 7/95] says of the Elmwood Ave. branch: “My wife and I eat there regualrly and really like it. There were some problems with consistency of quality for a while, but they seem to have straightened that out. Good service; good food; and nice, casual atmosphere.”

[NEW a+] That Coffee Place. UB Commons. 636–0494. Coffees, desserts, sandwiches, salads.

[REV **] Towne Restaurant. 186 Allen St. (in Allentown, east of Elmwood), Buffalo. 884–5128. Also: 827 Military Rd. (near Kenmore Ave.), Kenmore; 871–1092. Also: 120 Delaware Ave. (Lawley Bldg., downtown), Buffalo; 852–9750. (Recommendations are for the (original) Allentown branch only.) Greek, informal, inexpensive. Good soups, open souvlaki, spanakopita, baklava. The closest thing Buffalo has to a European café in ambiance; open 24 hrs. (but closes Sun at 11 P.M.). [J&AD, 1993] say “This plebeian Greek restaurant is a popular neighborhood hangout. Inexpensive, interesting clientele. Runyanesque atmosphere.” [RM, 5/94] says: “Pretty good Greek fare for lunches if you are in the Allentown area.”

[REV] Truffles. 1141 Kenmore Ave., Kenmore. 877–1111. Swiss. [AR, 9/93] says: “Went a couple of times before I went on sabbatical. Not very impressed.” One of [KNL]’s “top-ranked favorites” (5/95). Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.

[a+] Uni–Mart. UB Commons. Groceries, sandwiches, pizza.

[REV] Vito’s Market. 206 S. Elmwood Ave. (between Allentown and downtown), Buffalo. 852–5650. [J&AD, 1993] say: “Vito’s is a gourmet Italian Deli featuring unusual soups, sandwiches, salads, and delicious desserts, at reasonable prices.” [RM, 5/95] says: “Nice sandwiches—they are known for their hot sauces and have a huge selection.” Given 3 \frac{1}{2} (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[REV] Warren’s. 561 Main St., Tonawanda. 694–3700. [KNL] once said: “It has … excellent American cuisine in an elegant setting. I would put it in the same class as Rue Franklin,” but he adds (5/95) that he hasn’t been there in a while. Given 3 \frac{1}{2} (out of 4) stars by Janice Okun, Buffalo News restaurant reviewer.

[***] White Inn. 52 E. Main St. (Rt. 20), Fredonia. 672–2103. Excellent continental cuisine in a restored hotel (built in 1868).

[a] Woo Chon Korean Restaurant. 402 Evans Rd. (Pierce Arrow Plaza), Amherst. 626–5980. Includes “do-it-yourself Korean barbecue”. [PJE, 5/94] says: “Very good.”

[NEW a+] Young Chow. UB Commons. 689–6303, 6305. Chinese. They deliver.

[NEW a*] ZuZon American Grille. 5110 Main St. (just off the 290, in the Walker Center plaza), Williamsville. 634–6123. “California-style”. If you sit at, or near, the counter in the back, you get to watch the chefs: a free show with your meal. [RM, 7/95] says: “The latest Amherst Yuppie restaurant. Considering the effort they put into the atmosphere of the walls and general decor, one would think they would do something about the dreadful ceiling. The menu is equally uneven with no real theme, as are the patrons (anything from yuppie shorts to suits seems to go, and everyone from business people to young families seems represented). The food is also uneven. Some of the entrees were good, some of the side dishes quite poor. The desserts were well presented, including the now-standard plate painting, but mediocre at best. The one thing that was consistent was the high price, which presumably you are paying for the right to say you have been.” [AN&BS, 7/95] say: “Imaginative but soggy. Worth a try.” Given 4 (out of 5) stars by Janice Okun, Buffalo News restaurant reviewer.